OpenStudy (anonymous):

Please help me! Select the best possible first step to solving the system by first eliminating the x variable. 3x + 5y = 1 2x + 4y = –4 Multiply the first equation by 2 and multiply the second equation by 3. Multiply the first equation by –2 and multiply the second equation by 3. Multiply the first equation by –2 and multiply the second equation by –3. None of the above

OpenStudy (anonymous):

Multiply the first equation by –2 and multiply the second equation by 3.

OpenStudy (anonymous):

By doing above procedure the given eq will become -6x -10y = -2 6x + 12y = –12 ---------------- 2y =-14 (Adding both the equations) y=-7

OpenStudy (anonymous):

Thank you sooo much aain! :) Can I ask you two more questions?

OpenStudy (anonymous):

Oh sure..

OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 2y = 2 x + 3y = 17 (2y + 2) + 3y = 17 (–2y + 2) + 3y = 17 x + 3(2y + 2) = 17 x + 3(–2y + 2) = 17

OpenStudy (anonymous):

(2y + 2) + 3y = 17

OpenStudy (anonymous):

Thank you sooo much! Yhay's all for now!